OpenStudy (anonymous):

find the value of 10 - p for each given value of p 1.p=2.55 2.p=1.55 3.p=0.55 4.p=7.89 help :D

OpenStudy (anonymous):

The expression is 10 - p. Your task is to replace the "p" with each value in turn and then to work out the numerical answer. Look up "substitution" for extra help.

OpenStudy (anonymous):

1. 10-(2.55)=7.45 2.10-(1.55)=8.45 3.10-(0.55)=9.45 4.10-(7.89)=2.11 you just replace the p with the numerical values they give you in each problem